OpenStudy (anonymous):

Solve A = πr + 4t for t.

jimthompson5910 (jim_thompson5910):

The idea here is to isolate t. So we need to move that pr term to the left side (by subtracting it from both sides), then we need to divide both sides by 4 to completely isolate t. So this means.... \[\Large A = \pi r + 4t\] \[\Large A - \pi r = 4t\] \[\Large 4t = A - \pi r\] \[\Large t = \frac{A - \pi r}{4}\] So the solution is \[\Large t = \frac{A - \pi r}{4}\]

jimthompson5910 (jim_thompson5910):

sry made a typo, meant to write "πr" instead of "pr" up at the top, but everything else is ok.